Sven’s Component

Sven Nys is a divisional manager of Vandelay Industries. Sven is considering introducing a new component to the existing product line. Introducing the new component would require the purchase of new equipment. The equipment would cost $170,000, last for 2 years, and have zero salvage value at the end of the two years. Sven estimates that the variable overhead associated with the new component will be $95,000 a year and is allocated based on direct labor dollars. The new product will also require hiring a new employee at an annual wage of $30,000/year. This employee manufactures the new component. The new component consumes $20/unit of direct materials. The production volume, sales, and marketing costs for this new product line in each of the two years that the new component would be produced are as follows:

Year

2012

2013

Production (units)

1,800

1,700

Sales (units)

1,200

2,300

Selling Price/u

$150

$150

Marketing Cost

$10,000

$10,000

Vandelay Industries uses full absorption for both book and tax purposes. It is company policy that straight-line depreciation is used for all production machinery, and that fixed overhead is allocated on the basis of units produced. Vandelay Industries uses LIFO (last in, first out) inventory flow assumption. Vandelay Industries uses a 12% discount rate for all net present value calculations and faces a tax rate of 30% of net income. Van delay industries has positive income elsewhere in the business, so the taxes associated with any losses result in cash saved.

The company must pay for the machine in cash upon delivery. For purposes of the analysis, assume that that all other expenses associated with producing the new component will be paid in cash at the end of the year in which they occur. Also, assume that all sales of the new component will be received in cash at the end of the year in which they occur. Assume that all units produced in the year are complete and ending balance of WIP is 0.

Sven’s compensation is a base salary (regardless of wether the firm takes the project) plus a bonus of 10% of after-tax income, which is paid in cash at the end of the year. When considering the bonus scheme, the firm includes 30% of any loss as a reduction in the loss (e.g. a net income of -$100 results in an after-tax net income of -$70 and a "bonus" for SVEN of $-7). Please ignore the bonus when calculating the net income below and the net present value of the component from the firm's perspective below.

Question: What is the NPV of the project from Sven's perspective if production was equal to sales?
Answer #1

New Equipment Cost = $170000

So Cost of Initial Investment is $ 170000

  • Calculation Of Depreciation : Cost of Asset Value - Scrap Value => 170000 - 0   => 85000 Per Year

Life of an Asset Value 2

  • Calculation Annual Cash Flows :

Year 2012 Year 2013

Sales(Units) 1200 2300

Selling Price $150 $150

Revenue $180000 $345000

  • Calculation Of Present Value :

2012 2013

Annual Revenue 180000 345000

- Depreciation (85000) (85000)

Cash Flows After Dep 95000 260000

Tax @ 30% (28500) (78000)

Cash flows after tax 66500 182000

+ Depreciation 85000 85000

Cash Flows After Adj Dep   151500 267000

Discounted Rate @ 12% 0.892 0.797

Present Cash Flows 135138 212799

Total Cash Flows = 347937

Total Outflows = 170000 + 141828 + 144257 = 456085

NPV = 347937 - 456085 = - $108148

Note :

* Production is equal to Sales.